Circular No. 7636

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001cf IN UGC 7020 G. Sala, Azzate, Italy, reports the discovery with the 0.3-m reflector

at Promiod Observatory of an apparent supernova (mag 18.5) on unfiltered CCD images (limiting mag 19.5) taken on May 26.87 and 26.93 UT. The new object is located at α = 12^{h}02^{m}31^{s}.64, δ = +41^{o}02′58′′.9 (equinox 2000.0), which is 49′′.3 west and 19′′.1 south of the nucleus of UGC 7020. V. Pozzoli, Bareggio, Italy, confirmed the new object on unfiltered CCD images taken at the J. B. Postel Observatory on May 26.96, adding that both Palomar Sky Surveys show nothing at the position of SN 2001cf; he also adds that the new object is only ∼ 7′′ from a foreground star of red mag 16.9 in the USNO-A2.0 catalogue.

A. V. Filippenko and R. Chornock, University of California at Berkeley, report that inspection of a CCD spectrum (range 390–1000 nm), obtained on May 29 UT with the Keck-II 10-m reflector, reveals that SN 2001cf is a type-IIb supernova like SN 1993J (Filippenko et al. 1993, Ap.J. 415, L103). Hα absorption is strong, but lines of He I with P-Cyg profiles are also prominent. The recession speed of the host galaxy (from NED) is 6130 km/s.

SUPERNOVAE 2001ba, 2001bf, 2001bk, 2001bt M. Phillips, Carnegie Observatories; and K. Krisciunas, Cerro Tololo

Interamerican Observatory, report on spectra obtained with the Du Pont 2.5-m telescope at Las Campanas Observatory on May 17–18 and 26–27 UT. SN 2001bf (cf. IAUC 7620, 7625) is a type-Ic event, similar in appear-

ance to SN 2000H (cf. http://cfa-www.harvard.edu/cfa/oir/Research/

supernova/spectra/sn00h.gif). SN 2001bk (cf. IAUC 7624) is a type-II supernova at z ∼ 0.043 with strong, broad, P-Cyg Hα emission. SN 2001bt is a type-Ia event that was at or near maximum in a spectrum obtained on May 26.

Corrigenda. On IAUC 7614, SN 2001ba, the α offset should read 15′′.8 east. On IAUC 7633, SN 2001bt, the α offset should read 14′′.1 west.

SUPERNOVA 2001bs IN UGC 10018 A. V. Filippenko and R. Chornock, University of California at Berkeley,

report that a CCD spectrum (range 390–1000 nm), obtained on May 29 UT with the Keck-II 10-m reflector, reveals that SN 2001bs (cf. IAUC 7631) is a normal type-Ia supernova, ∼ 1 week past maximum brightness.

2001 May 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green